OpenStudy (anonymous):

Acylinder has a volume of 175 pi cubic units and a height of 7 units. The diameter of the cylinder is

Parth (parthkohli):

You may equate. First of all, the volume of a cylinder is \(\pi r ^2 h\)

OpenStudy (anonymous):

volume of cylinder =πr^2h diameter=2r pi=3.14 use them

Parth (parthkohli):

\[\pi r^2 h = 175 \Longrightarrow 3.14 \times r^2 \times 7 = 175 \] Clear?

Parth (parthkohli):

Solve for \(r\), then multiply that by 2. Remember, radius * 2 = diameter.

OpenStudy (anonymous):

those are not the answeres i have

OpenStudy (anonymous):

i have 5 units 10 units 12.5 units

Parth (parthkohli):

Solve the equation, please?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

just place h=7 in v=πr2h where v=175 and find r double val of r it gives diameter

OpenStudy (anonymous):

thanks]

OpenStudy (anonymous):

what is the answer because i still dont understand